# A Survey on State-of-the-art Deep Learning Applications and Challenges

### v5

@@ 6 Summary and Future Directions
Model Architecture: The recent advancements in computer vision focus on vision transformers (ViT) for scalable representation learning [18]. In image classification and image segmentation, ViTs are utilized to capture global context and relationships between different parts of an image, while in object detection, ViTs are used to streamline the detection process by eliminating the traditional components like anchor boxes and non-maximum suppression [carion_end--end_2020]. However, ViTs lack the ability to exploit local spatial features and struggle with hyperparameter sensitivity and performance on smaller datasets. To address this issue, researchers have developed hybrid deep learning architectures such as conformer [90] and MaxViT [91] that combine both transformers and convolutional neural networks to capture both local and global features. Similar trend is observed in modern natural language processing whereby the basis of the deep learning models is transformer architectures such as BERT [141] and its variants to achieve remarkable accuracy in tasks such as text classification, machine translation and text generation.
@@
While recent advancements in computer vision and natural language processing have predominantly relied on transformer architectures, the field of time series analysis focuses on a different trajectory. Notably, areas such as human activity recognition, speech recognition and financial prediction have seen significant progress through the development of hybrid architectures that do not depend on transformers. Instead, these advancements leverage recurrent neural networks and convolutional neural networks, often incorporating attention mechanisms. The hybrid architectures allow the models to capture both local features and temporal dependencies in the data, which is crucial for time series analysis [196, ige2023deep]. Similar trends can be observed in ECG and EEG classification whereby not only convolutional and recurrent neural networks are used, but also generative models such as variational autoencoder [268] and generative adversarial network to generate data to address challenges posed by limited datasets and overfitting [252]. The generator of the generative model is composed of layers of BiLSTM, while the discriminator is composed of convolutional neural networks.
@@
Contextual Enhancement: Recent advancements in text classification focus on capturing contextual information utilizing approaches such as joint embeddings of labels [148] and words as well as aspect-aware methods that enhance feature extraction [150]. Conversely, in machine translation, although transformers have proven effective, they often struggle to capture nuanced language features and contexts. To address this issue, various strategies have been introduced such as concatenating contextual sentences and context-sensitive training [158, 159]. In image generation, the researchers focus on improving the quality and relevance of generated images by enhancing the alignment between input text and images, while also addressing issues like fuzzy shapes and diversity through innovative deep learning architectures such as incorporating attention mechanisms and fusion modules into the model. Furthermore, recent models like DF-GAN [130] and DMF-GAN [131] showcase a shift towards single-stage generators with regularization strategies that maintain details while enhancing diversity. In text generation, BERT and its variants are the basis of the deep learning models. The research focuses on hierarchical architectures, capturing long text dependencies [171] and leveraging domain knowledge through label sequence [172] and conversation sequence [168] to bridge semantic gaps. Overall, the advancements show a clear trend toward improving contextual awareness and feature representation.
@@
Future Directions: Looking ahead, the future of deep learning presents numerous opportunities for growth and innovation. Future models could explore non-sequential hybrid architectures of transformer and convolutional neural networks to leverage the strengths of both approaches, enhancing performance in image classification, object detection and image segmentation. Furthermore, researchers could investigate simpler or lightweight architectures and new training schemes to address the long training time associated with transformers. Future research in image generation may explore advanced frameworks or architectures that further integrate semantic understanding, perhaps by employing hierarchical attention modules and/or fusion modules to capture both local and global features more effectively. Additionally, incorporating unsupervised learning or self-supervised learning approaches could reduce reliance on labeled datasets, allowing models to learn from a diverse range of inputs. In the area of natural language processing, future research in text classification and machine translation could explore the integration of external knowledge bases and domain-specific embeddings may further improve context understanding and label alignment. Additionally, refining context incorporation methods such as dynamic context updating during translation may enhance the model performance.
@@
Future research in image generation may explore advanced frameworks or architectures that further integrate semantic understanding, perhaps by employing hierarchical attention modules and/or fusion modules to capture both local and global features more effectively. Additionally, incorporating unsupervised learning or self-supervised learning approaches could reduce reliance on labeled datasets, allowing models to learn from a diverse range of inputs. In text generation, advancing towards more interactive and adaptive systems that can maintain context over extended dialogues and narratives is crucial. Furthermore, frameworks that can leverage multi-modalities (text, audio and visual) for richer contextual understanding may facilitate the development of more advanced applications across these domains. Future research can also focus on exploring adversarial attacks in these domains and developing tailored defence mechanisms. Also, researchers can further investigate the practical implications of adversarial attacks in real-world scenarios, such as in autonomous vehicles and medical imaging. Understanding the potential impact of adversarial attacks in these applications can inform the development of more robust and secure systems.
@@
Due to the sensitive nature of financial research, future work can focus on enhancing the interpretability of deep learning models in financial predictions. Researchers should explore techniques to explain the predictions of models, to improve trust and understanding of model decisions, which is essential for adoption in human activity recognition, speech recognition and finance. Furthermore, future models could explore the use of transformers to enhance feature representation for time series analysis. Future research can also incorporate attention mechanisms into hybrid generative model architectures to improve the quality of generated signals. Additionally, real-time processing capabilities are paramount, future work can develop efficient algorithms for real-time processing of ECG and EEG signals. This could involve optimizing existing architectures and leveraging hardware acceleration techniques to enable real-time inference on resource-constrained devices such as wearable sensors and implantable devices. Future models could explore methods for adapting or personalizing models to account for inter-subject variability and improve performance on individual subjects. Also, EEG electrodes cover only a fraction of the brain’s surface, resulting in limited coverage of neural activity. Deep learning models could investigate strategies to infer activity from unobserved brain regions or integrate information from multiple modalities to provide more comprehensive coverage. These areas can still be further explored.
@@ 7 Conclusion
Deep learning has become the prominent data-driven approach in various state-of-the-art applications. Its importance lies in its ability to revolutionize many aspects of research and industries and tackle complex problems which were once impossible to overcome. Numerous surveys have been published on deep learning, reviewing the concepts, model architectures and applications. However, the studies do not discuss the emerging trends in the state-of-the-art applications of deep learning and emphasize the important traits and elements in the models. This paper presents a structured and comprehensive survey of deep learning, focusing on the latest trends and advancements in state-of-the-art applications such as computer vision, natural language processing, time series analysis and pervasive computing. It explores key elements and traits in modern deep learning models, highlighting their significance in addressing complex challenges across diverse domains. The discussion also covers future research directions in advancing these fields. Furthermore, this paper presents a comprehensive review of the deep learning fundamentals, which is essential for understanding the core principles behind modern deep learning models. The survey finishes by discussing the critical challenges and future directions in deep learning.

### v6

@@ 6 Summary and Future Directions
Model Architecture: The recent advancements in computer vision focus on vision transformers (ViT) for scalable representation learning [21]. In image classification and image segmentation, ViTs are utilized to capture global context and relationships between different parts of an image, while in object detection, ViTs are used to streamline the detection process by eliminating the traditional components like anchor boxes and non-maximum suppression [126]. However, ViTs lack the ability to exploit local spatial features and struggle with hyperparameter sensitivity and performance on smaller datasets. To address this issue, researchers have developed hybrid deep learning architectures such as conformer [107] and MaxViT [108] that combine both transformers and convolutional neural networks to capture both local and global features. Similar trend is observed in modern natural language processing whereby the basis of the deep learning models is transformer architectures such as BERT [162] and its variants to achieve remarkable accuracy in tasks such as text classification, machine translation and text generation.
@@
While recent advancements in computer vision and natural language processing have predominantly relied on transformer architectures, the field of time series analysis focuses on a different trajectory, where transformers research is lagging due to fundamental challenges in data structure, sequence length and dataset size and availability. Notably, areas such as HAR and financial prediction have seen significant progress through the development of hybrid architectures that do not depend on transformers. Instead, these advancements leverage recurrent neural networks and convolutional neural networks, often incorporating attention mechanisms. The hybrid architectures allow the models to capture both local features and temporal dependencies in the data, which is crucial for time series analysis [222, 223]. The hybrid attention mechanisms improve the feature extraction process by dynamically weighting important channel, spatial and temporal features, enhancing model performance [224, 225, 226]. GAN models are used to address data scarcity in HAR, where the generator consists of convolutional layers [233, 234], with LSTM [231] while the discriminator consists of convolutional layers.
Similar trends can be observed in ECG and EEG classification, where convolutional and recurrent neural networks are commonly used to build hybrid models [299, 331], with attention mechanisms [303, 300]. Generative models such as variational autoencoder [325] and generative adversarial network are used to generate data to address challenges posed by limited datasets [308, 309, 337], poor quality [338], noisy signals [339], and overfitting [307]. The generative approach employs GAN only or with encoder-decoder [308], where the generator consists of convolutional layers with attention mechanisms or bidirectional LSTM, while the discriminator consists of convolutional layers with attention mechanisms.
@@
Contextual Enhancement: Recent advancements in text classification focus on capturing contextual information utilizing approaches such as joint embeddings of labels [169] and words as well as aspect-aware methods that enhance feature extraction [171]. Conversely, in machine translation, although transformers have proven effective, they often struggle to capture nuanced language features and contexts. To address this issue, various strategies have been introduced such as concatenating contextual sentences and context-sensitive training [179, 180]. In text generation, additional contexts are provided to improve both the quality and diversity of the generated sentences such as incorporating the category context [194] and style context [195] into text generation process.
In image generation, the researchers focus on improving the quality and relevance of generated images by enhancing the alignment between input text and images, while also addressing issues like fuzzy shapes and diversity through innovative deep learning architectures such as incorporating attention mechanisms and fusion modules into the model. Furthermore, recent models like DF-GAN [151] and DMF-GAN [152] showcase a shift towards single-stage generators with regularization strategies that maintain details while enhancing diversity. In text generation, BERT and its variants are the basis of the deep learning models. The research focuses on hierarchical architectures, capturing long text dependencies [193] and leveraging domain knowledge through label sequence [196] and conversation sequence [190] to bridge semantic gaps. Overall, the advancements show a clear trend toward improving contextual awareness and feature representation.
LLMs are increasingly being leveraged in robotics for improving human-robot interaction. The main challenge of integrating LLMs is bridging the gap between robots and the physical world to provide real-world contexts, a process known as LLM grounding. The concept of LLM grounding refers to how machines make sense of abstract language such as words and ideas by connecting them to objects they can actually experience or sense in the real world. To this end, sensors such as cameras and microphones, along with object detection, pose estimation and speech recognition methods, are used for semantic and perceptual grounding. The robot data are streamed are at varying sampling rates, so they need to be aligned and carefully selected to identify key events [355]. The detected objects and their relationships are then modeled using 3D scene graph for scene understanding [353]. Central to the grounding system is a module that collates and constructs the text prompts which are sent to the LLM [354]. The text prompts are typically formulated by integrating a set of predefined robot skills that can be executed [352].
@@
Future Directions: Looking ahead, the future of deep learning presents numerous opportunities for growth and innovation. Future models could explore non-sequential hybrid architectures of transformer and convolutional neural networks to leverage the strengths of both approaches, enhancing performance in image classification, object detection and image segmentation. Furthermore, researchers could investigate simpler or lightweight architectures and new training schemes to address the long training time associated with transformers. Future research in image generation may explore advanced frameworks or architectures that further integrate semantic understanding, perhaps by employing hierarchical attention modules and/or fusion modules to capture both local and global features more effectively. Additionally, incorporating unsupervised learning or self-supervised learning approaches could reduce reliance on labeled datasets, allowing models to learn from a diverse range of inputs.
@@
In the area of natural language processing, future research in text classification and machine translation could explore the integration of external knowledge bases and domain-specific embeddings may further improve context understanding and label alignment. Additionally, refining context incorporation methods such as dynamic context updating during translation may enhance the model performance. In text generation, advancing towards more interactive and adaptive systems that can maintain context over extended dialogues and narratives is crucial. Furthermore, frameworks that can leverage multi-modalities (text, audio, and visual) for richer contextual understanding may facilitate the development of more advanced applications across these domains. Future research can also focus on exploring adversarial attacks in these domains and developing tailored defense mechanisms. Furthermore, researchers can further investigate the practical implications of adversarial attacks in real-world scenarios, such as in autonomous vehicles and medical imaging. Understanding the potential impact of adversarial attacks in these applications can inform the development of more robust and secure systems.
@@
Real-time processing capabilities are paramount, future work can develop efficient models or algorithms for real-time processing of HAR, speech recognition, ECG and EEG signals. This could involve optimizing existing architectures and leveraging hardware acceleration techniques to enable real-time inference on resource-constrained devices such as wearable sensors and implantable devices. Furthermore, future models could explore lightweight transformers to enhance feature representation for time series analysis. Due to the sensitive nature of financial research, future work can focus on enhancing the interpretability of deep learning models in financial predictions. Researchers should explore techniques to explain the predictions of models, to improve trust and understanding of model decisions, which is essential for adoption in HAR, speech recognition and finance.
Future research in data generation for time series analysis and pervasive computing should focus on developing more advanced techniques to address data scarcity and improve model generalization. In HAR, frameworks that can generate diverse signals for subjects of different ages, genders, as well as physical abilities, may facilitate the development of models that can generalize across different populations and real-world scenarios [405]. Moreover, the generated data should capture sensor noise, variations in user behavior, and diverse environmental conditions. In finance, synthetic market data should model complex dependencies, volatility patterns and anomalies to improve trend and risk predictions. For EEG and ECG classification, researchers could explore deep generative models that can create physiologically meaningful synthetic signals while preserving individual variability. Future models could explore methods for adapting or personalizing models to account for inter-subject variability and improve performance on individual subjects. Furthermore, EEG electrodes cover only a fraction of the brain’s surface, resulting in limited coverage of neural activity. Deep learning models could investigate strategies to infer activity from unobserved brain regions or integrate information from multiple modalities to provide more comprehensive coverage. These areas can still be further explored.
Future research in robot object identification and navigation will focus on enhancing real-time perception, generalization across diverse environments, and increasing robustness to occlusions and dynamic obstacles. Furthermore, researchers can explore multimodal learning, integrating vision with LiDAR and tactile sensing to improve scene understanding. Advances in self-supervised learning and few-shot learning could enable robots to identify and navigate complex environments with minimal labelled data. LLMs will play a crucial role in human-robot interactions, facilitating natural language understanding, intent recognition and context-aware decision-making [406]. Researchers should optimize LLMs for real-time processing on robots and improve their grounding in physical environments. Additionally, future work can explore a deep learning-based navigation with LLMs to develop more autonomous, assistive and socially intelligent robots.
The future of AI lies beyond deep learning, moving towards a conscious intelligent system that mimics human cognition [407]. Achieving this requires a significant advancement in hardware and software architectures, enabling systems to process information with greater efficiency, adaptability and reasoning capabilities. Beyond mere pattern recognition, future of AI will need to integrate symbolic reasoning, causal inference and common sense to truly emulate human thought [408, 409, 410]. Future research will likely focus on hybrid AI models that combine deep learning with symbolic AI, enabling machines to understand abstract concepts, reason about cause and effect and apply knowledge across domains. Quantum computing, with its unique property of entanglement could exponentially accelerate deep learning training and inference [411, 412]. Future models could explore quantum neural networks which could process high dimensional data more effectively and discover patterns that are not possible by classical methods. Quantum-based optimizers could also solve problems that are infeasible for classical graphical processing units or tensor processing units. A more radical approach is neuromorphic computing that uses specialized hardware to mimic the structure and function of the human brain, enabling low-power inference, parallel processing and real-time learning and adaptation [413, 414]. This results in autonomous AI systems that can solve complex tasks requiring quick response time and adapt to changing situations. Finally, multi-modal intelligence enables AI to process and integrate multiple types of sensory input, including vision, speech and touch, providing a more holistic view of the world, similar to human perception [415].
@@ 7 Conclusion
Deep learning has become the prominent data-driven approach in various state-of-the-art applications. Its importance lies in its ability to revolutionize many aspects of research and industries and tackle complex problems which were once impossible to overcome. Numerous surveys have been published on deep learning, reviewing the concepts, model architectures and applications. However, the studies do not discuss the emerging trends in the state-of-the-art applications of deep learning and emphasize the important traits and elements in the models. This paper presents a structured and comprehensive survey of deep learning, focusing on the latest trends and advancements in state-of-the-art applications such as computer vision, natural language processing, time series analysis and pervasive computing, and robotics. It explores key elements and traits in modern deep learning models, highlighting their significance in addressing complex challenges across diverse domains. The discussion also covers future research directions in advancing these fields. Furthermore, this paper presents a comprehensive review of the deep learning fundamentals, which is essential for understanding the core principles behind modern deep learning models. The survey finishes by discussing the critical challenges and future directions in deep learning.